Grade 12Algebra

If a,b,c are in hp and a^2,b^2,c^2 are in hp then ?

Nikhil
9 Years agoGrade 12

1 Answer

Shahid
9 Years ago
if a,b,c are in H.P , then
 2/b = 1/a +1/c                                                        --------------{1}
       = (a+c)/ac
2ac = ab+bc 
ac = (ab+bc)/2                                                          --------------{2}
 
And Given a2 , b2 , c are in H.P.
Therefore , 2/b2  = 1/a2 +1/c2
                         = (a2+c2)/a2c2
2a2c2 = a2b2+b2c2
a2c= (a2b2 + b2c2)/2                                               -------------{3}
 
{(ab+bc)/2}2 = (a2b2+b2c2)/2
(a2b2 + b2c2 + 2ab2c)/4   = (a2b2 + b2c2)/2
2ab2c = a2b2 + b2c2 
(a2b2 + b2c2) – 2(ab)(bc) = 0 
(ab-bc)2 = 0
ab= bc
a =c                                                               ------------{4}
 
From {1} , 
2/b = 1/a+ 1/c
      = 1/a +1/a
      = 2/a
a=b                                                                ----------------{5}
 
From {4} , {5} , we get,
a=b=c.